OpenStudy (anonymous):

the resultant of two forces 3N and 4N making an angle 90 with each other is a) 1N b) 7N c) 5N d) 3.5N

OpenStudy (anonymous):

the answer would be 5N apply this formula for resultant \[\sqrt{A ^{2}+B ^{2}+2AB \cos \theta}\] cos90=0 so, \[\sqrt{4^{2}+3^{2}}=\sqrt{16+9}=\sqrt{25}=5\]

OpenStudy (anonymous):

is it correct equation?

OpenStudy (anonymous):

they are at 90 degree with each other. why is not sin theta in place of cos theta?

OpenStudy (anonymous):

i am not quite sure about answer!you can ask someone who is confirmed about it.

OpenStudy (anonymous):

you can take it as forula,if you ever want to calculate the resultant of two vectors,this formula is used